# Robustness Optimization for Compact Free-electron Laser Driven by Laser Wakefield Accelerators

### v4

@@ Abstract
Despite the successful demonstration of compact free electron lasers (FELs) driven by laser wakefield accelerators (LWFAs), the shot-to-shot fluctuations inherent to LWFAs remain a major obstacle to realizing LWFA-driven FELs with high gain and robust operation. Here, we present a conceptual design for LWFA-driven FELs with enhanced robustness and reliability. By employing Bayesian optimization techniques, the beamline was optimized to achieve sufficient tolerance against these fluctuations under actual experimental conditions. Start-to-end simulations revealed that this systematic optimization significantly reduces the system’s sensitivity to parametric variations. With the optimized configurations, the radiation energy can be maintained above 1 μJ at a wavelength of approximately 25 nm, even when accounting for twice the root-mean-square (RMS) ranges of these inherent jitters. This proposed scheme represents a substantial advancement in the development of compact LWFA-driven FEL systems, enabling robust operation and paving the way for the realization of reliable and widely accessible sources.
@@
Free electron lasers (FELs) are capable of generating high-brilliance and coherent radiation across a broad range of wavelengths, extending even into the hard X-ray regime [1, 2, 3, 4, 5]. Their unique capabilities hold the potential to revolutionize various scientific disciplines by enabling unparalleled atomic resolutions on femto-to-attosecond timescales [6, 7]. These FEL facilities typically rely on the radio-frequency (RF) accelerators with large scales and incurs exorbitant costs. With unprecedented accelerating gradients reaching several hundreds of gigavolt per meter [8, 9], several orders of magnitude higher than those of RF accelerators, laser wakefield accelerators (LWFAs) provide a promising alternative for driving compact FELs. The realization of such compact LWFA-driven FELs has been identified as one of the major challenges in this decade, as addressed in the European Plasma Research Accelerator with eXcellence In Applications (EuPRAXIA) [10]. Motivated by the vast potential of these more accessible and cost-effective systems, numerous innovative schemes and state-of-the-art technologies have been proposed, resulting in significant improvements in beam qualities of the accelerated e beams from LWFAs [11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Recently, compact FELs driven by LWFAs have been experimentally demonstrated, with the self-amplified spontaneous emission (SASE) configuration at 27 nm and seeded configuration at 269 nm [21, 22]. However, the reliability and reproducibility issues of LWFAs present major obstacle for realizing compact and robust FELs for application [23, 24].
@@
In LWFAs, the interaction between high-intensity laser pulses and plasma constitutes an exceptionally complex and nonlinear physical process. Minor fluctuations of the experimental conditions, particularly in the drive laser and plasma properties, can cause substantial shot-to-shot fluctuations in e-beam qualities [11]. Furthermore, these fluctuations in beam qualities, in terms of energy, energy spread and emittance, can lead to varying degrees of beam property degradation during transport through the beamline. These shot-to-shot fluctuations hinder the gain and pose significant robustness challenges for LWFA-driven FELs [23]. To address these challenges, it is imperative to systematically quantify the beam quality fluctuations induced by the inherent jitter of both laser and plasma parameters. Additionally, substantial optimization of the beamline is essential to enhance its tolerance to these shot-to-shot fluctuations, thereby ensuring robustness of LWFA-driven FELs under actual experimental conditions. Bayesian optimization (BO) has emerged as a particularly suitable approach for optimizing such complex nonlinear system with high evaluation costs or prolonged evaluation times, and has been demonstrated as an effective optimization tool in the field of laser-plasma physics and accelerators [25, 26, 27, 28, 29, 30].
@@
In this Letter, we present a conceptual design for high-gain LWFA-driven FELs with enhanced robustness and reliability, which is crucial for practical applications. Through detailed numerical simulations, the shot-to-shot fluctuations in beam quality have been quantitatively analyzed, which stem from both laser and plasma parameters, including laser energy variation, focal position displacement induced by wavefront distortion, and shock front position instability in the plasma. To address these fluctuations, BO techniques were employed for beamline optimization, utilizing the power gain length and FEL energy as the objective function. Start-to-end simulations show that the optimized system can maintain FEL radiation in the saturated or near-saturated regime with corresponding energy exceeding 1 μJ in the extreme ultraviolet (EUV) region, even operating within twice the root-mean-square (RMS) range of each parameter jitter. In addition, we conducted a quantitative analysis of the system’s tolerance to e beam pointing jitter, revealing that the proposed scheme can tolerate pointing jitters up to 1 mrad while maintaining FEL radiation energy above 1 μJ. This proposed scheme represents a significant advancement in the development of compact LWFA-driven FEL systems with robust operation, offering a promising pathway toward the realization of reliable and widely accessible system.
@@
To investigate the shot-to-shot fluctuations in beam quality generated by LWFAs, quasi-three-dimensional (3D) particle-in-cell (PIC) simulations were conducted using the FBPIC code with a co-moving simulation box [31, 32]. The simulation utilized a Gaussian laser pulse with a central wavelength of $\lambda=800$ nm, a waist radius of $\omega_{0}=35$ μm, a pulse duration of 25 fs (full width at half maximum, FWHM), and a normalized vector potential of $a_{0}$ = 1.3. The plasma density profile was designed based on the previously observed fringe pattern, and the synergistic mechanism of the injection was implemented, as detailed in Ref. [33]. Figure 1(a) illustrates the density profile used in the simulation, with the shock front length configured to 75 μm.
@@
Figure 1: PIC simulation results of LWFA with inherent parameter jitters obtained from experiments. (a) Schematic of plasma density profile and fluctuation sources, with the pink shaded area region and the orange solid line represent the range of focal position jitter and the initial shock-front position. (b)-(g) Electron beam properties with inherent parameter jitters: the blue, red and orange curves represent the laser energy variations, laser focal position displacement and shock front position instability, respectively. Horizontal axes denotes ±$\pm 2$ folds RMS range of these three parameter jitters, and the vertical axes represent beam energy (b), charge (c), slice energy spread (d), global energy speread (e), horizontal (f) and vertical (g) normalized projected emittance.
@@
The shot-to-shot fluctuations of the accelerated e beams primarily originate from inherent jitters in both the laser and plasma. These intrinsic jitters can be quantitatively described by three specific parameters: laser energy variation, focal position displacement caused by wavefront distortion, and shock front position jitter in the plasma. Through systematic experimental measurements, the corresponding RMS values were determined to be $\Delta E/E=\pm 0.54\%$ for relative laser energy variation [34] and $\sigma_{zshock}=4.9$ μm for shock front position jitter, respectively. While the focal position displacement induced by wavefront distortion was not directly measured in our experiments, we adopted a reported value of $\sigma_{zfoc}=0.2$ mm according to Ref. [17]. The details of the measured shock front jitter were presented in Appendix AA. A comprehensive quantitative characterization of the e-beam parameters under the influence of these inherent laser and plasma jitters is presented in Figs. 1(b)- 1(g), including beam energy, charge, slice energy spread, global energy spread, as well as horizontal and vertical normalized projected emittance. It is noted that the laser energy variation is represented by the variation of laser normalized vector potential $a_{0}$ in the following discussions. These jitters span ±$\pm 2$ times the measured RMS values in the presented simulations. Each e-beam parameter depicted in Figs. 1(b)- 1(g) containing 13 independent simulation runs, with $\Delta_{RMS}=0$ serving as the initial and jitter-free condition.
@@
In the presented simulations, a synergistic mechanism was employed, wherein both the shock front and the self-evolution of the laser contributed to the injection process. By precisely controlling electron injection and chirp compensation through tailored plasma density profiles, high-quality electron beams can be obtained [33]. However, inherent parameter jitter introduces significant instability in electron injection, leading to substantial charge fluctuations. As shown in Fig. 1(c), when accounting for jitter in all three parameters ($\sigma_{a_{0}}$, $\sigma_{zfoc}$, and $\sigma_{zshock}$), the relative charge fluctuation reaches ±$\pm 22.0\%$ rms. Such pronounced charge variations induce distinct beamloading effects across different jitter conditions, resulting in variations in acceleration efficiency and inhomogeneities in the accelerating fields. Consequently, the beam quality degrades markedly, with relative energy fluctuation of ±$\pm 2.7\%$ rms, slice energy spread fluctuation of ±$\pm 21.2\%$ rms, global energy spread fluctuation of ±$\pm 17.3\%$ rms, horizontal normalized projected emittance fluctuations of ±$\pm 16.2\%$ rms and vertical normalized projected emittance fluctuations of ±$\pm 15.1\%$ rms.
@@
These instabilities fundamentally limit the radiation gain and stability of FELs, where achieving high-gain operation typically necessitates three requirements: elevated charge, reduced slice energy spread, and minimized transverse emittance [23]. Our analysis reveals an intrinsic incompatibility among these three metrics, as evidenced by their anti-correlated trends in Figs. 1(c),  1(d) and  1(g). The phase space manipulation of electron beams via beamline optics enables precise control and optimization of beam quality parameters. By leveraging this fundamental principle, strategic design and optimization of the beamline architecture can effectively mitigate beam quality discrepancies, thereby establishing a balanced parameter space essential for generating stable high-gain FEL radiation. This approach addresses a critical factor in maintaining the consistency of FEL performance by compensating for inherent variations in electron beam characteristics through phase space manipulation.
@@
$q_{6}$20 0.0 -11.55
Figure 2: The optimization results of the $E_{min}$ for the 13-electron ensemble. (a) The simulated objective function $E_{min}$ (dots) with the cumulative best results (black dashed line). (b)-(c) The evolution of the beam size (b) and the normalized projected emittance (c) along the beamline in horizontal (blue) and vertical (orange) directions. The tracking beamline was configured using the best parameters obtained from the optimization.
To achieve high-gain LWFA-driven FELs with reliable and robust operations, it is essential to design a beamline with sufficient tolerance to accommodate the shot-to-shot fluctuations of the e beam induced by experimental jitters, as discussed in Fig. 1. The beamline consists of the following components: three permanent magnetic quadrupoles with a length of 5 cm each, three electromagnetic quadrupoles with a length of 10 cm each, and a 180-period planar undulator with a period length of 25 mm and an undulator parameter of $K_{0}=1.41$. These parameters was chosen according to our experimental conditions as reported in Ref. [21, 33]. A BO approach was employed for the beamline optimization. The optimization parameters included the strength parameters of the quadruples and the distances between each element, arranged in sequence from upstream to downstream as follows: $l_{0}$, $q_{1}$, $l_{1}$, $q_{2}$, $l_{2}$, $q_{3}$, $l_{3}$, $q_{4}$, $l_{4}$, $q_{5}$, $l_{5}$, $q_{6}$, $l_{6}$. Here, $l_{0}$ represents the distance between the gas target and the first quadruple, $l_{1}-l_{5}$ denote the distances between adjacent quadruples, and $l_{6}$ corresponds to the distance between the last quadrupole and the undulator entrance. The variables $q_{1}-q_{6}$ represent the strength parameters of the quadruples.
Figure 3: Radiation energy as a function of (a) normalized vector potential $\Delta a_{0}$, (b) laser focal position $\Delta{z_{foc}}$ and (c) the shock-front position $\Delta{z_{shock}}$. Mean values (black squares) and statistical uncertainties (standard deviation) calculated from 20 independent simulations with varying random seeds, where the blue curve represents optimization using $<L_{G}^{-1}>$ as objective function and the red curve represents optimization using $E_{min}$ as the objective function.
Figure 4: Simulation results for FELs with the 13-electron-beam ensemble. (a) Radiation energy along the undulator. (b) Pulse duration and (c) spectrum of the generated radiation pulses at the exit of the undulator. The red solid line represents jitter-free, and the red transparents line represent different jitters. Each line represents the average value of 20 distinct simulations with random seeds.
@@
We employ the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [35] implemented within the Optuna framework [36] as the primary optimization methodology. Within the context of BO, the formulation of the objective function is of curial importance, as it serves as the mathematical embodiment of our target optimization metric. Here, two distinct objective functions were implemented: the mean reciprocal value of the 3D power gain length $<L_{G}^{-1}>$ and the minimization of the simulated radiation energy $E_{min}$ obtained from the 13-electron-beam ensemble shown in Fig. 1. Both objective functions were strategically optimized within a systematic framework to achieve maximization of the desired performance metrics. Before the optimization of FEL radiation energy, we used the $<L_{G}^{-1}>$ optimization to narrow down the parameters of the beamline components. This constrained parameter space serves preventing the particle loss event during FEL simulations and enhancing the computational reliability of the optimization process. The details of the $<L_{G}^{-1}>$ optimization were presented in Appendix BB.
@@
In the optimization framework targeting $E_{min}$, we conducted parallel particle tracking simulations utilizing the 13-electron-beam ensemble illustrated in Fig. 1. The results of the beam tracking obtained from ELEGANT code [37] served as the input for subsequent FEL simulation performed with 3D and time-dependent code GENESIS [38]. The parameter space for this optimization, detailed in Table 1, was systematically refined based on the results from $<L_{G}^{-1}>$ optimization.
Figure 2(a) illustrates the evolution of the objective function $E_{min}$ as a function of iterations during the BO process. It should be noted that the radiation energy was obtained by a single run with specific seed for speed-up propose during the optimization process. The optimal solution achieves at iteration 11,020, with the corresponding optimal parameters detailed in Table 1. Under these optimized conditions, the radiation energy consistently exceeds 1 μJ for all electron beams in the ensemble, indicating robust operation of the LWFA-driven FELs in the saturation or near saturation regime and effectively accommodating shot-to-shot fluctuations. The corresponding evolution of the beam sizes and transverse emittance along the beamline is presented in Figs. 2(b) and  2(c), respectively. Figures 3(a)- 3(c) provide the comparison of the two optimization frameworks targeting $<L_{G}^{-1}>$ and $E_{min}$ under their respective optimized conditions. It is noted that the FEL energy was simulated with distinct random seeds for 20 runs in each situation in Fig. 3. While both frameworks exhibit similar trends in radiation energy, the $E_{min}$ optimized framework demonstrates superior performance with higher radiation energy. This enhancement can be attributed to two aspects, the fitting distortion of Eq. (B) for the short-pulse situation induced by the slippage effects [39], and the potential deviations introduced by using slice-averaged parameters for power gain length calculations. Through numerical simulations, the slice beam properties can be optimized to simultaneously achieve localized small averaged beam size, energy spread, and high beam current, where the lasing occurs. Such optimization ensures more efficient electron-radiation interaction and facilitates energy extraction efficiency from the e beam.
Figure 4 presents the properties of the FEL radiation with the optimal operation targeting $E_{min}$. The evolution of the radiation along the undulator for the 13-electron-beam ensemble is shown in Fig. 4(a). A minority population enters the exponential gain regime near the 1-meter position of the undulator, while the majority initiates exponential growth beyond 1.5 meters, approaching saturation or approximate saturation at end of the undulator. Figure 4(b) illustrates the radiation power, indicating localized lasing due to the precisely designed beamline parameters that impact localized phase-space characteristics through energy spread minimization and emittance matching, thereby enabling enhanced gain amplification dynamics. Notably, even the minimum radiation power maintains sub-GW level operation, demonstrating remarkable tolerance to system fluctuations. The corresponding radiation wavelength fluctuates with the RMS value of 0.8 nm, causing by the energy fluctuations of the accelerated e-beam, and thus the resonant wavelength, as illustrated in Fig. 4(c). It should be noted that each line in Fig. 4 denotes the average value over 20 independent simulations with random seeds.
Beam pointing stability is another critical issue that must be addressed to ensure the robust operation of LWFA-driven FELs. To quantitatively evaluate the pointing instability, artificial angular kicks were introduced at the entrance of the beamline. Using e beams simulated under jitter-free conditions ($\Delta_{RMS}=0$), we artificially introduced angular deviations in both transverse planes ($x^{\prime}:0.1-1.0$ mrad, $y^{\prime}:0.1-1.0$ mrad). Figure 5 presents the quantitative analysis of radiation energy degradation caused by this angular perturbation range, revealing a monotonic decrease in output energy from 34 μJ in the jitter-free case to 1 μJ with a pointing deviation of 1 mrad. Even with a pointing deviation of 1 mrad, the average radiation energy, simulated over 20 random seeds, remains above 1 μJ. These results indicate that the proposed scheme exhibits remarkable tolerance to angular deviations.

### v5

@@ Abstract
Despite the successful demonstration of compact free electron lasers (FELs) driven by laser wakefield accelerators (LWFAs), the inherent shot-to-shot fluctuations in LWFAs, including both laser and plasma instabilities, remain a primary obstacle to realizing LWFA-driven FELs with robust operation. Here, we present a conceptual design for LWFA-driven FELs with sufficient tolerance against shot-to-shot fluctuations using the Covariance Matrix Adaptation Evolution Strategy (CMA-ES). Start-to-end simulations demonstrated that this systematic optimization resulted in a significant improvement in the robustness of FELs. With the optimized configurations, the radiation energy can be maintained above 1 µJ at a wavelength of approximately 25 nm, even when accounting for twice the root-mean-square (RMS) ranges of these instabilities. This proposed scheme represents a substantial advancement in the development of compact LWFA-driven FEL systems, enabling robust operation and paving the way for the realization of reliable and widely accessible sources.
I Introduction
@@
Free electron lasers (FELs) are capable of generating high-brilliance and coherent radiation across a broad range of wavelengths, extending even into the hard X-ray regime [33, 13, 23, 28, 9]. Their unique capabilities hold the potential to revolutionize various scientific disciplines by enabling unparalleled atomic resolutions on femto-to-attosecond timescales [4, 7]. These FEL facilities typically rely on the radio-frequency (RF) accelerators with large scales and incurs highly costs. With unprecedented accelerating gradients reaching several hundreds of gigavolt per meter [38, 5], several orders of magnitude higher than those of RF accelerators, laser wakefield accelerators (LWFAs) provide a promising alternative for driving compact FELs. The realization of such compact LWFA-driven FELs has been identified as one of the major challenges in this decade, as addressed in the European Plasma Research Accelerator with eXcellence In Applications (EuPRAXIA) [2]. Motivated by the vast potential of these more accessible and cost-effective systems, numerous innovative schemes and state-of-the-art technologies have been proposed, resulting in significant improvements in beam qualities of the accelerated e beams from LWFAs [14, 18, 8, 6, 39, 11, 34, 29, 16, 19]. Recently, compact FELs driven by LWFAs have been experimentally demonstrated, with the self-amplified spontaneous emission (SASE) configuration at 27 nm and seeded configuration at 269 nm [40, 31]. However, the limited reliability and reproducibility of LWFAs present a fundamental barrier to realize compact and operationally robust FELs for practical application [21, 17].
@@
In LWFAs, the interaction between high-intensity laser pulses and plasma constitutes an extraordinarily complex, nonlinear process. Consequently, even modest fluctuations in the drive laser or plasma parameters, inevitably result in substantial shot-to-shot fluctuations in e-beam qualities [14]. Moreover, these fluctuations in beam qualities, including energy, energy spread and emittance, degrade the beam during transport through the beamline, ultimately hinder the gain and pose significant robustness challenges for LWFA-driven FELs [21]. To address these challenges, it is imperative to systematically quantify the beam quality fluctuations induced by the instabilities of both laser and plasma. Additionally, substantial optimization of the beamline is essential to enhance its tolerance to these shot-to-shot fluctuations, thereby ensuring robustness of LWFA-driven FELs. For optimizing such complex nonlinear systems with high evaluation costs, black-box optimizers have emerged as a particularly suitable approach. Various optimizers, including Bayesian optimization, genetic algorithms, and evolutionary strategies like Covariance Matrix Adaptation Evolution Strategy (CMA-ES), have been demonstrated as effective tools in the field of laser-plasma physics and accelerators [12, 25, 30, 10, 22, Döpp2023].
@@
In this paper, we propose a conceptual design for a high-gain LWFA-driven FELs with enhanced robustness and reliability. Through detailed numerical simulations, the shot-to-shot fluctuations in beam quality have been quantitatively analyzed, which are attributable to instabilities from both laser and plasma, including laser energy jitter, focal position displacement induced by wavefront distortion, and shock front position instability in the plasma. To address these fluctuations, we employed the CMA-ES algorithm for beamline optimization, with the corresponding objective function of the FEL energy. Start-to-end simulations show that the optimized system can maintain FEL radiation in the saturated or high-gain regime with corresponding energy exceeding 1 µJ in the extreme ultraviolet (EUV) region, even operating within twice the root-mean-square (RMS) range of each parameter jitter. Quantitative tolerance analysis further demonstrates the scheme’s robustness to up to 1 mrad of beam pointing jitter while maintaining radiation energy above 1 µJ. This energy corresponds to more than $1\times 10^{11}$ photons per pulse in EUV regime, satisfying the photon flux requirements for applications such as coherent diffractive imaging [37, 27]. As a conceptual design with clear pathways for future experiment, the proposed scheme represents a significant step toward robust, compact LWFA-driven FELs and offers a promising route to reliable and widely accessible systems.
II LWFA SIMULATION
@@
To investigate the shot-to-shot fluctuations in beam quality generated by LWFAs, quasi-three-dimensional (3D) particle-in-cell (PIC) simulations were conducted using the FBPIC code with a co-moving simulation box [32, 24]. The simulation domain has a size of 50 µm and 120 µm in the longitudinal and transverse directions, respectively, with the longitudinal grid size $\Delta z=31.25$ nm and the transverse grid size $\Delta r=80$ nm and 16 macroparticles per cell. In the simulation, we used a Gaussian laser pulse with a central wavelength of $\lambda=800$ nm, a waist radius of $\omega_{0}=35$ µm, a pulse duration of 25 fs (full width at half maximum, FWHM), and a normalized vector potential of $a_{0}$ = 1.3. The plasma density profile employed in the simulation is shown in Fig. 1(a), with a peak shock density of $3.9\times 10^{18}$ cm-3, a plateau density of $2.2\times 10^{18}$ cm-3, and a shock width of 75 µm [15].
@@
Figure 1: PIC simulation results of LWFA with inherent parameter jitters. (a) Schematic of plasma density profile and fluctuation sources, with the pink shaded area region and the orange solid line represent the range of focal position jitter and the initial shock-front position. (b)-(g) Electron beam properties with inherent parameter jitters: the blue, red and orange curves represent the laser energy variations, laser focal position displacement and shock front position instability, respectively. Horizontal axes denotes $\pm 2$ folds RMS range of these three parameter jitters, and the vertical axes represent beam energy (b), charge (c), slice energy spread (d), global energy speread (e), horizontal (f) and vertical (g) normalized projected emittance.
@@
The shot-to-shot fluctuations of the accelerated e beams primarily originate from instabilities from both laser and plasma. These instabilities were quantitatively described laser energy jitter, focal position displacement induced by wavefront distortion, and shock front position instability in the plasma. In the presented simulation, the relative RMS laser energy jitter of $\Delta E/E=\pm 0.54\%$ [41] is applied, corresponding to the performance of the 200 TW Ti:sapphire laser system operating at 1 Hz [41]. The focal position displacement jitter is set to an RMS value of $\sigma_{zfoc}=0.2$ mm, based on the reported values [34]. For the shock front position, an RMS jitter of $\sigma_{zshock}=4.9$ µm is adopted, following the experimental data provided in Appendix A. A comprehensive quantitative characterization of the e-beam parameters under the influence of these instabilities of laser and plasma is presented in Figs. 1(b)- 1(g), including beam energy, charge, slice energy spread, global energy spread, as well as horizontal and vertical normalized projected emittance. It is noted that the laser energy jitter is represented by the variation of normalized vector potential $a_{0}$ in the following discussions. These jitters were varied up to ±2 times the RMS values in the simulations. Each e-beam parameter depicted in Figs. 1(b)- 1(g) containing 13 independent simulation runs, with $\Delta_{RMS}=0$ serving as the initial and jitter-free condition.
@@
By precisely controlling electron injection and chirp compensation through tailored plasma density profiles, high-quality electron beams can be obtained [15]. However, these parameter jitters destabilizes the electron injection process, leading to substantial charge fluctuations. As shown in Fig. 1(c), when accounting for jitters in all these three parameters ($\sigma_{a_{0}}$, $\sigma_{zfoc}$, and $\sigma_{zshock}$), the relative charge fluctuation reaches $\pm 22.0\%$ RMS. Such charge variations lead to distinct beamloading effects under different jitter conditions, which in turn cause changes in acceleration efficiency and inhomogeneities in the accelerating fields. Consequently, the beam quality degrades markedly, exhibiting the following RMS fluctuations: $\pm 2.7\%$ in energy, $\pm 21.2\%$ in slice energy spread, $\pm 17.3\%$ in global energy spread, $\pm 16.2\%$ and $\pm 15.1\%$ in horizontal and vertical projected emittance, respectively.
@@
These instabilities fundamentally limit the radiation gain and operational stability of FELs, for which high-gain lasing generally demands three key beam properties: elevated charge, reduced slice energy spread, and minimized transverse emittance [21]. Our analysis reveals an intrinsic trade-off among these three metrics, as shows by their anti-correlated trends in Figs. 1(c),  1(d) and  1(g). Strategic design and optimization of the beamline layout can effectively mitigate beam quality discrepancies, thereby defining a workable parameter space that supports stable, high-gain FEL radiation. This approach compensates for the inherent fluctuations in electron beam characteristics, directly addressing a critical challenge in achieving consistent FEL performance.
III BEAMLINE OPTIMIZATION
Achieving reliable, high-gain LWFA-driven free-electron lasers (FELs) requires a beamline design robust enough to accommodate the shot-to-shot fluctuations of the e beam caused by experimental jitters, as discussed in Fig. 1. The beamline consists of three 5-cm-long permanent-magnet quadrupoles, three 10-cm-long electromagnetic quadrupoles, and a 180-period planar undulator with a period length of 25 mm and an undulator parameter of $K_{0}=1.41$. The layout of the beamline is depicted in Fig. 2.
Figure 2: Schematic of the optimized beamline.
@@ III BEAMLINE OPTIMIZATION
$q_{6}$ -20 0.0 -11.55
@@
We employ CMA-ES [20] implemented within the Optuna framework [1] as the primary optimization methodology. With the default setting, we adopt the theoretical optimum population size formulation $\lambda=4+3\ln n$ and standard adaptation rules for covariance matrix and step-size updates as established in the CMA-ES literature. The algorithm was initialized with a step size of 0.5 relative to normalized parameter bounds. The formulation of the objective function is crucial to the optimization process, as it serves as the mathematical embodiment of the target metric. Here, two distinct objective functions were implemented: the mean reciprocal value of the 3D power gain length $L_{G}^{-1}>$ and the minimization of the simulated radiation energy $E_{min}$ obtained from the 13-electron-beam ensemble shown in Fig. 1. Both objective functions were strategically optimized within a systematic framework to achieve maximization of the desired performance metrics. The optimization parameters included the strength parameters of the quadruples and the distances between each element, arranged in sequence from upstream to downstream as follows: $l_{0}$, $q_{1}$, $l_{1}$, $q_{2}$, $l_{2}$, $q_{3}$, $l_{3}$, $q_{4}$, $l_{4}$, $q_{5}$, $l_{5}$, $q_{6}$, $l_{6}$. Here, $l_{0}$ represents the distance between the gas target and the first quadruple, $l_{1}-l_{5}$ denote the distances between adjacent quadruples, and $l_{6}$ corresponds to the distance between the last quadrupole and the undulator entrance. The variables $q_{1}-q_{6}$ represent the strength parameters of the quadruples. To accelerate the subsequent FEL energy optimization, the reciprocal of the gain length $L_{G}^{-1}>$ was first maximized to identify a constrained set of viable beamline parameters. The detail of the $L_{G}^{-1}>$ optimization were presented in Appendix B.
Figure 3: The optimization results of the $E_{min}$ for the 13-electron ensemble. (a) The simulated objective function $E_{min}$ (dots) with the cumulative best results (black dashed line). The y-axis is plotted on a logarithmic scale. (b)-(c) The evolution of the beam size (b) and the normalized projected emittance (c) along the beamline in horizontal (blue) and vertical (orange) directions. The tracking beamline was configured using the best parameters obtained from the optimization.
@@
While minimizing the gain length $L_{G}$ establishes a beamline conducive to efficient FEL amplification, it can lead to distortion for FEL gain evaluation, as $L_{G}$ is evaluated from slice-averaged values of the beam. To address this, we shift the optimization objective to the minimum FEL energy $E_{min}$ expected within the 13-electron-beam ensemble show in Fig. 1. The results of the beam tracking obtained from ELEGANT code [3] served as the input for subsequent FEL simulation performed with 3D and time-dependent code GENESIS [36]. The parameter space for this optimization, detailed in Table 1, was systematically refined based on the results from $L_{G}^{-1}>$ optimization.